RC meeting and Run Plan, Thursday September 22nd
From SBS wiki
RC Meeting Information
Run Coordinator: Donald Jones
Location: Counting House Conference Room
Counting House status
Subsystem status reports
- Hall A Status -- Jesse Butler
- Target status/timeline (training?) -- Arun
- Beam line:
- Monitors (calibration) -- Simona
- Correctors -- Bogdan, Iuliia
- Wiki, shift schedule, shift checklist -- Don
- Tentative run plan document -- Bogdan
- Analyzer -- Andrew
- DAQ -- Alex
- HCal -- Jiwan
- Spectrometer magnets -- Jack
- GEMs/Tracking -- Holly
- Are new HV controls, alarms, gas flow monitoring procedures, EPCIS tracking in place?
- Shift checklist ok?
- Shower Calorimeter -- Provakar
- Beam polarization -- Don
- RC report:
- Laser safety committee was not able to approve the laser for operation due to an inactive critical safety component, the smoke sensor in the laser room. Some interlocks were not working as expected but were successfully fix by Ellen in real time.
- Schedule has been adjusted to reflect a Tuesday start, but might need further cancelations. Need people to sign up for shifts.
- Shower calorimeter: Provakar reminded us that for each new setting of the magnet currents and positions, the shower calorimeter needs to be recalibrated due to magnetic fields affecting PMTs and thus the trigger efficiency. He will provide a run plan for this part of the program.
- GEMs: Holly reported that the GEMs are essentially ready. The SBS GEMs take a mix of gas that will require daily changeouts of tanks. She is working with procurement to ensure that there is an automatic weekly supply of these tanks delivered and on a scheduled changeout plan (expert driven). She has provided a run plan for commissioning the GEMs including a luminosity scan to check the new HV voltage dividers and a HV scan to measure efficiency. These need to be done early on and will be expert-driven.
- Beamline: Simona is working with Yves to perfect the process for establishing beam in the hall. Yves had mentioned perhaps touching the edge of the collimator with beam to see the contrast for IC calibration. Bogdan strongly protested ANY touching of the copper collimator with beam (not even tune); Silviu said it would be ok from a target standpoint. Simona reminded us that the "empty target" really means threading the beam through a 1" gap between the reference cell and the optics targets and is not out of beam like we generally think. We need to exercise care because of this.
- HCal: Jiwan told us that HCal is ready but will need some commissioning time to ensure the timing is good and to get the coincidence trigger online.
- DAQ: no update, but Bob Michaels says it is being used routinely.
- Analyzer: Andrew reports that the analyzer is ready for operation although there are some things that still need to be tested related to the 3 stream setup.